Help! Please help with this.My teacher did not explain how to go about this.
"write an equation to describe the exponential function as y =ab^x. The base is
4. The graph passes through the point (3,5)"

Thank you so much for the help!

Answer by scott8148(6628)

you are given b (the base) and the point (3,5) or (x,y)

plug in the values to find a ___ 5 = a * 4^3 ___ 5 = a * 64 ___ 5/64 = a

y = (5/64) 4^x